OpenStudy (anonymous):

What is the measure of each base angle of an isosceles triangle if its vertex angle measures 38 degrees and its 2 congruent sides measure 21 units?

OpenStudy (campbell_st):

draw a diagram  so usding angle sum of a triangle 180 = 38 + x + x or 180 = 38 + 2x all you need to do is solve for x